OpenStudy (anonymous):

Factor Completely: 5x2 - 40

OpenStudy (jackellyn):

5(x^2-8)

OpenStudy (anonymous):

Thanks!

OpenStudy (jackellyn):

No problem :)

OpenStudy (unklerhaukus):

oh but you can go one step further \[5x^2 - 40\]\[=5(x^2-8)\]\[=5(x^2-(2\sqrt2)^2)\]\[=5(x+2\sqrt2)(x-2\sqrt 2)\]

OpenStudy (anonymous):

Okay, thanks!